What is the direct labor rate variance? 50 unfavorable 125 unfavorable 125 favorable The following information for Q 7-8 The St. Augustine Corporation originally budgeted for $360,000 of fixed overhead at 100% normal production capacity. Production was budgeted to be 12,000 units. The standard hours for production were 5 hours per unit. The variable overhead rate was S3 per hour. Actual fixed overhead was $360,000 and actual variable overhead was $170,000. Actual production was 11,800 units. 7. The variable factory overhead controllable variance is 7,000 favorable $7,000 unfavorable c)$6,000 favorable $6,000 unfavorable 8. The fixed factory overhead volume variance is $7,000 favorable 7,000 unfavorable $6,000 favorable $6,000 unfavorable 9. Morocco Desk Co. purchases 6,000 feet of lumber at $6.00 per foot. The standard price for direct materials is $5.00. The entry to record the purchase and unfavorable direct materials price variance is
Answer #1

7.

Actual variable overhead = $170,000

Budgeted variable overhead rate per unit = Budgeted hours per unit x Budgeted variable overhead rate per hour

= 5 x 3

= $15

Actual output = 11,800 units

Budgeted variable overhead for 11,800 units = 11,800 x 15

= $177,000

Variable factory overhead controllable variance = Actual variable overhead - Budgeted variable overhead

= 170,000 - 177,000

= $7,000 Favorable

Correct option is (A)

8.

Budgeted fixed overhead for 12,000 units = $360,000

Fixed overhead rate per unit = 360,000/12,000

= $30  

Absorbed fixed overhead = Actual output x Standard fixed overhead rate per unit

= 11,800 x 30

= $354,000

Budgeted fixed overhead = $360,000

Fixed factory overhead volume variance = Absorbed fixed overhead - Budgeted fixed overhead

= 354,000 - 360,000

= $6,000 favorable

Correct option is (c)